The average starting salaries for 55 graduates with a certain major was $43949. The standard deviation for all graduates with that major was $8937. Give the 94% confidence interval for the population mean. Round the endpoints to the nearest dollar, and do not put commas in your numbers.

Confidence interval: ( , )
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 43949

sample standard deviation = s = 8937

sample size = n = 55

Degrees of freedom = df = n - 1 = 54

At 94% confidence level the z is ,

\alpha = 1 - 94% = 1 - 0.94= 0.06

\alpha / 2 = 0.06 / 2 = 0.03

t\alpha /2,df = t0.03,54 = 1.921

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.921 * ( 8937/ \sqrt 55)

= 2314.930

The 99% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

43949 - 2314.930 < \mu < 43949 + 2314.930

41634.07 < \mu < 46263.93

(41634 , 46264)